A clock battery wears out after moving 87780 C of charge through the clock at a rate of 42 mA. How many days before the batteries run out?


Round your answer to 2 decimal places


1
Expert's answer
2021-05-10T19:11:32-0400

We can find the number of days before the battery runs out from the formula:


"Q=It,""t=\dfrac{Q}{I},""t=\dfrac{87780\ C}{42\cdot10^{-3}\ A}=2.09\cdot10^6\ s\cdot\dfrac{1\ h}{3600\ s}\cdot\dfrac{1\ day}{24\ h}=24\ days."